OpenStudy (anonymous):

Absolute Value of y=3 (solution or no solution) Absolute value of n + 2 = 4(solution or no solution)

OpenStudy (anonymous):

For the first one, the absolute value of what number will give you 3?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

yes. so there is a solution for that equation. agree?

OpenStudy (anonymous):

By the way, is there another number, when you take the absolute value of that number, you will get 3?

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

No, absolute value of 0 is 0. What is the absolute value of -3?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay. Can you help me with a few more problem that i'm stuck on please?

OpenStudy (anonymous):

what about your second problem?

OpenStudy (anonymous):

Absolute value of n+2 is 4. Tell me, the absolute value of what number will give you 4?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

You mean, n = 2?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

correct, when n= 2, you have n+2 = 4

OpenStudy (anonymous):

Is there another value of n that will work?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I think there is. Besides 4, what other number can you think of, when you take the absolute value, will give you 4?

OpenStudy (anonymous):

I don't know.

OpenStudy (anonymous):

what about -4?

OpenStudy (anonymous):

ohh yeahh

OpenStudy (anonymous):

Isant the absolute value of -4...4?

OpenStudy (anonymous):

Isn't*

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so take n+2 and set it equal to -4. and you can solve that equation...and what do you get for n?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

no...to solve n+ 2 = -4 subtract 2 from both sides, you get n = what?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

no, -4-2 = -6

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so your second equation does have solutions. That answers that question.

OpenStudy (anonymous):

so now you have solved your two questions. They both have solutions.

OpenStudy (anonymous):

okay. thank you.

OpenStudy (anonymous):

welcome.